Walter White “found” $1,000,000 and deposited it into his checking account at Wells Fargo Bank. Answer the following questions in response to this new $1,000,000 in checkable deposits. The reserve ratio is 8 percent.

a. What is the max. amount of new loans Wells Fargo Bank can make (i.e. excess reserves, E)?

b. What is the size of the monetary multiplier, m?

c. Calculate maximum checkable-deposit creation, D.

d. Answer questions a,b,c on the assumption that the reserve ratio is 5 percent.

e. In response to the economic shutdown caused by the COVID-19 pandemic, The Federal Reserve's new corporate-credit facility took in $305 million worth of exchange-traded funds on the first day of its operation, according to data released Thursday

▪ Briefly explain the cause-effect chain from this specific expansionary monetary policy (i.e. the Fed buying corporate-debt ETFs), and how the Fed expects this action to boost real GDP in the U.S. economy. You can simply view this as a form of quantitative easing.
Answer #1

a) Excess reserve = Checkable deposits - required reserve

                              = 1,000,000 - (1,000,000 * 0.08)

                              = $920,000

Thus, the max. amount of new loans Wells Fargo Bank can make is $920,000.

b) The size of the monetary multiplier = 1 / Required reserve ratio = 1 / 0.08 = 12.5

c) The maximum checkable-deposit creation = 1,000,000 * 12.5 = $12,500,000.

d) If the reserve ratio is 5 percent,

Excess reserve = Checkable deposits - required reserve

                              = 1,000,000 - (1,000,000 * 0.05)

                              = $950,000

Thus, the max. amount of new loans Wells Fargo Bank can make is $920,000.

The size of the monetary multiplier = 1 / Required reserve ratio = 1 / 0.05 = 20

The maximum checkable-deposit creation = 1,000,000 * 20 = $20,000,000.